∠A is an acute angle in a right triangle. Given that sinA=725, what is the ratio for cosA? Enter your answer in the boxes as a fraction in simplest form
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52778)

.
∠A is an acute angle in a right triangle. Given that sinA=725, what is the ratio for cosA?
Enter your answer in the boxes as a fraction in simplest form
~~~~~~~~~~~~~~~~~~

Sin(A) can not be equal to 725.


///////////////////         \\\\\\\\\\\\\\\\\\\\\\


                    To the managers of this project  ( ! )


As  I  notice,  nonsensical messages arrive to this forum,  created by persons illiterated in  Math.


Examples are these posts

https://www.algebra.com/algebra/homework/logarithm/logarithm.faq.question.1192804.html

https://www.algebra.com/algebra/homework/Linear-equations/Linear-equations.faq.question.1192802.html

https://www.algebra.com/algebra/homework/Pythagorean-theorem/Pythagorean-theorem.faq.question.1192801.html

https://www.algebra.com/algebra/homework/logarithm/logarithm.faq.question.1192800.html



I ask you please to take all necessary measures and remove the person/(persons) who created these posts
from their positions in the project due to their inability to do their job properly.




Answer by Alan3354(69443)

∠A is an acute angle in a right triangle. Given that sinA=725, what is the ratio for cosA? Enter your answer in the boxes as a fraction in simplest form
===================
There is a solution for sin(A) = 725, but it's a complex number.
I doubt that's what you meant to ask.
========================
I suspect it's sin(A) = 7/25
Then cos(A) = sqrt(1 - sin^2) = sqrt(1 - 49/625) = sqrt(576/625)
cos(A) = 24/25
---------------------------------------------------
Don't post problems with typos.
Look before you submit.